OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21

OpenStudy (itiaax):

To find the y intercept, substitute 0 for x. Therefore: 0+7y=-21 y=-21/7 y-intercept = -3 To find the x-intercept, substiture 0 for y. Therefore: x+7(0)=-21 x=-21

OpenStudy (anonymous):

X-Int: -21 Y-Int: -3

OpenStudy (anonymous):

so could it be written like this: x intercept = (-3, 0); y intercept = (0, -21)

OpenStudy (itiaax):

Yes!